Decide whether (-4,2) is a solution of the system.
3x+2y=-8
2x-y=-6

Answer by ewatrrr(24785)


Hi
Is (-4,2) a solution for this system: NO
3x+2y=-8 | -12 + 4 = -8
2x-y=-6 | -8 - 2 ≠ -6
2x-y=-6 OR y = 2x+6
3x+2y=-8
3x + 2(2x+6) = -8
7x = -20
x = -20/7 or -2 6/7 and y = 2/7 (-40/7 + 42/7)
CHECKING our Answer***
2*-20/7 - 2/7 = -42/7 = -6